a first order process has an experimentally determined rate constant of 6.0 x 10^-4 at 500 C if the initial conc of the species is 0.0226 mol/L calc it conc after 955 secs
Answer #1


first order has the integrated law equation is as follows

ln[A]t = -Kt + ln[A]o

where [A]o = initial concentration = 0.0226 M

K = 6.0*10-4

t = 955 sec

[A]t = concentration at time

substituting the values in the equation

ln[A]t = - 6.0*10-4 * 955 + ln [0.0226 ]

ln[A]t = - 6.0*10-4 * 955 + (-3.7898)

ln[A]t = - 6.0*10-4 * 955 -3.7898

[A]t = 0.01274 M

answer => 0.0127 M